Suppose utility function is given by U(x,y) = x0.3y0.7 subject to budget constraint I= Pxx + Pyy. Further suppose that the prices for those goods given by Px=1 Py=2 & I=2. Derive demand function of good x and good y
1
Expert's answer
2020-07-09T13:52:40-0400

"\frac {\delta U}{\delta x}=0.3(\frac{y}{x})^{0.7} = \lambda"

"\frac {\delta U}{\delta y}=0.7(\frac{x}{y})^{0.3}=2 \lambda"

"0.7(\frac{x}{y})^{0.3}=2\times 0.3(\frac{y}{x})^{0.7}"

"\frac {x}{y}=\frac {6}{7}"


"x=\frac {6}{7} I"


"1\times x+2\times y=I"

"\frac {6}{7}y+2y=I"

The demand function of good x and good y


"x=0.3I"

"y=0.35I"